OpenStudy (anonymous):

The perimeter of a square is 32 feet. What is the area of the square?

OpenStudy (anonymous):

P = s + s + s 32 = 4s S = 8 so : \[A = s^2= 8^2 = 64 ft^2\]

OpenStudy (anonymous):

P = s+s+s+s* ^^"

OpenStudy (anonymous):

Sweet... 64ft was a possibility

OpenStudy (anonymous):

awesome ^_^

OpenStudy (anonymous):

I have like 2 more questions left..

OpenStudy (anonymous):

let's see if we can do it!

OpenStudy (anonymous):

I asked another one over there.. lol thanks so much for helping me

OpenStudy (anonymous):

np ^_^